Let the production function of a firm is given as
1
Expert's answer
2021-10-11T16:46:17-0400

Solution:

a.). Derive MRTS = "\frac{MP_{L} }{MP_{K}}"

Let x = L

      y = K

MPL = "\frac{\partial Q} {\partial L} = \frac{1} {x^{0.5} }"


MPK = "\frac{\partial Q} {\partial K} = \frac{1} {y^{0.5} }"


"\frac{MP_{L} }{MP_{K}} = \frac{w }{r}"


"\frac{\frac{1} {x^{0.5} } }{\frac{1} {y^{0.5} } } = \frac{w} {r }"


"\frac{y^{0.5}} {x^{0.5} } = \frac{w} {r }"


x = "\frac{r^{2}y} {w^{2} }"


y = "\frac{w^{2}x} {r^{2} }"

 

b.). The cost function of the firm:

TC = rY + wX

TC = r("\frac{w^{2}x} {r^{2} }") + w("\frac{r^{2}y} {w^{2} }")

TC = "\frac{w^{2}x} {r} +" "\frac{r^{2}y} {w}"